At what frequency would the wavelength of sound in air be equal to the mean free path of oxygen molecules at 1.3 atm pressure and -2.0°C? Take the diameter of an oxygen molecule to be 5.7 × 10-8 cm and the speed of sound to be 320 m/s.
Answer #1

Mean free path ART a Erd²NAP 8 - 10 R = gas constant 8.314 J/mol-k T= (273- 2) = 271K d = = 5.7x10 cm = 5.7 X10 23 NA Avogado